### v2

# WeatherDiffusion: Controllable Weather Editing in Intrinsic Space

Abstract
We present WeatherDiffusion, a diffusion-based framework for controllable weather editing in intrinsic space. Our framework includes two components based on diffusion priors: an inverse renderer that estimates material properties, scene geometry, and lighting as intrinsic maps from an input image, and a forward renderer that utilizes these geometry and material maps along with a text prompt that describes specific weather conditions to generate a final image. The intrinsic maps enhance controllability compared to traditional pixel-space editing approaches. We propose an intrinsic map-aware attention mechanism that improves spatial correspondence and decomposition quality in large outdoor scenes. For forward rendering, we leverage CLIP-space interpolation of weather prompts to achieve fine-grained weather control. We also introduce a synthetic and a real-world dataset, containing 38k and 18k images under various weather conditions, each with intrinsic map annotations. WeatherDiffusion outperforms state-of-the-art pixel-space editing approaches, weather restoration methods, and rendering-based methods, showing promise for downstream tasks such as autonomous driving, enhancing the robustness of detection and segmentation in challenging weather scenarios.
1 Introduction
For autonomous vehicles, robust scene understanding requires the ability to handle adverse weather conditions. Techniques that handle weather effects, can substantially improve the robustness of perception models across different weather conditions. While diffusion-based image editing techniques [labs2025flux1kontextflowmatching, wu2025qwenimagetechnicalreport, brooks2023instructpix2pix, hertz2022prompttopromptimageeditingcross] have created significant opportunities here, a key limitation remains: the lack of fine-grained controllability in the generated scenarios.
Following existing image editing approaches, weather editing methods [cosne2020using, li2021weatherganmultidomainweather, tasnim2025normalizing] typically leverage a unified generative model to perform weather transformations in the pixel space. Particularly, WeatherWeaver [lin2025controllable] decomposes weather editing into a two-stage strategy: weather removal and weather synthesis. While these methods have achieved impressive results in weather manipulation, they still struggle to preserve the underlying material and geometry of the scene while generating natural illumination.
In this paper, we present WeatherDiffusion, a novel framework for controllable weather editing that operates in intrinsic space. Our key insight is that by decomposing a scene into its fundamental components—material properties, geometry, and lighting, we can achieve higher control over weather effects than in the image space. WeatherDiffusion consists of two key components: a forward renderer that uses CLIP-space interpolation and diffusion priors for fine-grained weather synthesis, and a novel inverse renderer tailored for outdoor and autonomous driving scenes. We are inspired by recent diffusion-based intrinsic decomposition and recomposition approaches, RGBX [zeng2024rgb] and DiffusionRenderer [liang2025diffusionrenderer], which have achieved impressive results at the indoor and object levels; however, they do not focus on weather and do not generalize to large-scale autonomous driving scenarios. We bridge this gap with a novel intrinsic map-aware attention mechanism that ensures spatial correspondence, thereby significantly enhancing decomposition fidelity in complex, unconstrained outdoor environments. By combining the inverse and forward renderers, our framework allows for controlled editing of weather and lighting.
Our WeatherDiffusion outperforms existing state-of-the-art inverse and forward rendering methods, achieving over 10 dB PSNR improvement in inverse rendering and higher PickScore [kirstain2023pickapicopendatasetuser] in forward rendering. Compared with weather restoration and pixel-space editing methods, our approach achieves higher text-image consistency and better DINO-based structural alignment, enabling cleaner and more controllable weather editing. By proactively correcting environmental distortions at the visual input level, we significantly boost the performance of downstream tasks such as object detection and semantic segmentation. We observe that after applying WeatherDiffusion, the detection and segmentation performance on the ACDC [sakaridis2021acdc] benchmark increases by 87.1% ( from 13.15% to 24.60%) and 24.5% ( from 24.13% to 30.05%), respectively. We also introduce a synthetic dataset and a real-world dataset containing 38k and 18k images with intrinsic map annotations, covering diverse weather conditions and a wide range of driving scenes to train these two components. To summarize, our contributions are as follows:
We propose WeatherDiffusion, a method to decompose images into intrinsic maps under various weather conditions, and synthesize them into another lighting or weather condition guided by text prompts.
We introduce intrinsic map-aware attention that provides customized visual detail guidance for generative models, helping our decomposition.
We construct two new datasets called WeatherSynthetic and WeatherReal, containing synthetic and real-world images covering various weather conditions on autonomous driving scenarios, along with their corresponding maps. The datasets will be released upon acceptance.
2 Related work
Diffusion models. Diffusion models have achieved remarkable progress in high-fidelity and text-conditioned image generation [ho2020denoising, rombach2022high, ramesh2022hierarchical, esser2024scaling]. Modern diffusion frameworks typically adopt a denoising process [ho2020denoising, song2020denoising, liu2022pseudo, lipman2022flow], parameterized by UNet- or DiT-based backbones [ronneberger2015unet, peebles2023dit]. In our research, we repurpose diffusion models to jointly estimate material, geometry, and lighting from an image while synthesizing new images under specified weather conditions. This demonstrates that the strong priors embedded in pre-trained diffusion models can be effectively leveraged for faithful and physically grounded estimations.
Forward and inverse rendering using diffusion.
The emergence of the diffusion model has catalyzed a novel methodology that leverages generative models to learn the joint probability distribution between images and their corresponding intrinsic maps [kocsis2024intrinsic, zeng2024rgb, li2024idarb, fu2024geowizard, chen2024uni, liang2025diffusionrenderer, he2025unirelightlearningjointdecomposition, IntrinsicDiffusion]. IID [kocsis2024intrinsic] focuses on material estimation in indoor scenes, while RGBX [zeng2024rgb] extends diffusion to bidirectional mapping between RGB images and maps. Several works further incorporate geometric priors [fu2024geowizard] or multi-view cues [li2024idarb]. DiffusionRenderer [liang2025diffusionrenderer] adapts a video diffusion model to achieve temporally consistent inverse and forward rendering, and UniRelight [he2025unirelightlearningjointdecomposition] jointly estimates albedo and relighted video frames. These methods are primarily designed for indoor scenes, small objects, or video relighting, and they struggle to generalize to large-scale outdoor driving scenes with diverse weather conditions. Moreover, none of these works address controllable weather editing or decomposition in intrinsic space under multiple weather conditions, which is the goal of our framework.
Image and weather editing.
Most image editing techniques operate purely in the pixel space, using diffusion models to modify color, texture, or local appearance without modeling underlying scene factors [tumanyan2022plugandplaydiffusionfeaturestextdriven, brooks2023instructpix2pix, wu2025qwenimagetechnicalreport, labs2025flux1kontextflowmatching, hertz2022prompttopromptimageeditingcross, cao2023masactrltuningfreemutualselfattention, parmar2023zeroshotimagetoimagetranslation, kawar2023imagictextbasedrealimage]. Following this paradigm, existing weather-editing models [cosne2020using, li2021weatherganmultidomainweather, tasnim2025normalizing] directly translate one weather type to another, and WeatherWeaver [lin2025controllable] finetunes a video diffusion model for weather removal and synthesis. However, pixel-space editing lacks physical interpretability and cannot guarantee consistent material, geometry, or illumination. Intrinsic-space manipulation has been explored by IntrinsicEdit [Lyu_2025], but it focuses on object-level editing rather than large-scale outdoor scenes. Other approaches edit weather in 3D space [qian2025weathereditcontrollableweatherediting, li2023climatenerfextremeweathersynthesis], but they require accurate geometry, which is unavailable for real-world driving scenes. In contrast, WeatherDiffusion performs controllable weather editing in the intrinsic space. This formulation enables fine-grained and geometry-preserving control that is difficult to achieve in the pixel space.
3 Method
3.1 WeatherDiffusion
We aim to modify weather-related factors (e.g., weather particles and accumulations) of a scene while preserving its underlying geometry and material properties. Pixel-space editing methods inherently entangle weather effects with appearance, making it difficult to maintain structural consistency and natural illumination across different weather conditions. Moreover, weather phenomena are primarily tied to a scene’s lighting, rather than intrinsic material attributes. These observations motivate us to move beyond direct pixel manipulation and instead operate in intrinsic space.
We propose WeatherDiffusion, a framework designed for controllable weather editing in intrinsic space. The framework includes two components: the weather-aware inverse renderer and the weather-conditioned forward renderer. The image is input into the inverse renderer, which performs intrinsic decomposition, disentangling images into weather-invariant material and geometry, as well as weather-variant illumination. Correspondingly, the forward renderer re-renders images based on given intrinsic maps and text prompts that specify the target weather or lighting.
We repurpose Stable Diffusion 3.5 [sd3.5_medium] to enable the inverse and forward renderers. To leverage the diffusion prior to achieve fine-grained weather control, we first obtain a weather transitional direction in the CLIP space:
where denotes the text prompt describing a specific weather type, and is the CLIP text encoder. Then we shift a weather-neutral embedding by steps along this direction:
Replacing the original prompt with can force the model to generate reasonable intermediate results. To preserve the rich priors of the pre-trained model, we align part of the forward renderer’s intermediate features with those of the original Stable Diffusion. The detailed implementation of feature distillation is shown in the supplementary material.
3.2 Intrinsic map-aware attention
Outdoor and autonomous driving scenes exhibit larger variations in object scale. As shown in Fig. 3, the original Stable Diffusion lacks explicit attention guidance and often performs poorly on distant, small objects and geometrically complex regions. We observe that different intrinsic maps require attention to distinct regions of an image, as shown in Fig. 4. To leverage these observations, we extend DiT with intrinsic map-aware attention (IMAA) to apply attention guidance for DiT.
We identify important regions based on the intrinsic maps and use them to construct an attention bias. Specifically, we first employ DINOv2 [oquab2023dinov2] to extract a set of patch tokens . For each intrinsic map, we define a learnable embedding that captures its inherent characteristics. A gating mechanism is applied to selectively filter patch tokens based on the current intrinsic map. Formally, we compute a map-aware mask:
where and are linear projections of patch tokens and map embedding , indicates the concatenated input to the MLP. This gating mechanism highlights image regions most relevant to the target map.
We construct a joint attention bias using to guide the diffusion model, effectively enhancing the attention logits in DiT. We apply the map-aware mask to the text–image and image–image parts of the joint attention matrix: the former enforces textual guidance on important regions, while the latter strengthens feature aggregation within the image space. Formally, the bias is defined as:
where and denote token indices, and is the set of image tokens. Then the bias is applied to the DiT:
where are combination of image and text token.
Moreover, we devise a heuristic-guided progressive training strategy for IMAA to stabilize learning and ensure that IMAA provides meaningful guidance in the early stages. For example, we use the gradient operator to extract illuminated regions and shadow boundaries for the irradiance map. The detailed description and ablation studies are shown in the supplementary material.
3.3 Dataset construction
Existing datasets of images with corresponding intrinsic maps suffer from the absence of outdoor environments (OpenRooms [li2020inverse], Hypersim [roberts2021hypersim], InteriorVerse [zhu2022learning]) or insufficient weather diversity (MatrixCity [li2023matrixcity]), and thus are inappropriate for large-scale outdoor and autonomous driving scenes. While MatrixCity provides continuous variations of fog density and illumination, it mainly targets relighting and reconstruction tasks. In contrast, our work requires diverse weather and lighting conditions rather than smooth transitions of the same type. Moreover, MatrixCity suffers from legal licensing issues. To fill the gap in large-scale weather-diverse autonomous driving datasets with paired images and intrinsic maps, we propose WeatherSynthetic and WeatherReal. Sample images from our datasets are shown in Fig. 5.
WeatherSynthetic
is a large-scale synthetic dataset encompassing a wide range of scene and weather types:
Weather: sunny, overcast, rainy, thunderstorm, snowy, foggy, sandstorm.
Time of day: early morning, morning, noon, afternoon.
Environment: urban, suburban, highway, parking
We use Unreal Engine 5 to render all images and intrinsic maps. We purchased 3D assets that are cleared for generative model use in Fab. The rendering pipeline uses the movie render queue and multi-sample anti-aliasing, producing high-quality rendering results. The UltraDynamicSky and UltraDynamicWeather are applied to modify the weather and daytime. Note that all images are in linear space without tone mapping. In total, rendering the 38K images and maps took about 24 hours on our setup.
WeatherReal
is a real-world dataset on autonomous driving scenes with various weather conditions. We use our inverse renderer to generate intrinsic maps of open-source datasets such as Waymo [sun2020scalability] and Kitti [geiger2013vision]. We use a multimodal model to remove challenging scenarios (e.g., rainy nights) to ensure the generation of high-quality pseudo-labels, and check the quality manually. The samples are shown in the supplementary materials. Moreover, we employ a pre-trained image editing model (i.e., Flux-Kontext [labs2025flux1kontextflowmatching]) to alter the weather types. Our WeatherReal is motivated by the observation that after training our model merely on synthetic data, the inverse and forward renderers lack sufficient generalization capability on real-world samples. Note that WeatherReal is only used to finetune models and is not used to evaluate results.
4 Experimental results
In this section, we first compare our method with pixel-space editing and weather restoration methods. Then we evaluate the inverse and forward rendering. We then conduct ablation studies on IMAA and datasets, and conclude with a discussion of the limitations of our approach.
Following WeatherWeaver [lin2025controllable], we use PickScore [kirstain2023pickapicopendatasetuser], CLIP image-text consistency (denoted as CLIP-S), and DINO structure similarity (denoted as DINO-S) to evaluate editing results. Following previous works [zeng2024rgb, li2024idarb], we report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Mean Angular Error (MAE), and Learned Perceptual Image Patch Similarity (LPIPS) for inverse rendering. We compare our performance with pixel-space editing methods (Flux-Kontext [labs2025flux1kontextflowmatching], Qwen-Image-Edit [wu2025qwenimagetechnicalreport], Instruct-Pix2Pix [brooks2023instructpix2pix], WeatherWeaver [lin2025controllable]) and weather restoration methods (AWRaCL [rajagopalan2025awracle], Histoformer [sun2024restoringimagesadverseweather]). We also compare our inverse and forward rendering results with RGBX [zeng2024rgb], IID [kocsis2024intrinsic], Geowizard [fu2024geowizard], IDArb [li2024idarb] and DiffusionRenderer [liang2025diffusionrenderer]. We evaluate different methods on WeatherSynthetic, Waymo [sun2020scalability], TransWeather [valanarasu2021transweather], ACDC [sakaridis2021acdc], and additional Internet images covering diverse weather conditions.
4.1 Comparison with pixel-space editing methods
We show quantitative comparisons in Tab. 1. Our method achieves the highest CLIP-S, indicating that it produces the most text-aligned and plausible weather editing results. In terms of DINO-S, we rank second only to Flux-Kontext [labs2025flux1kontextflowmatching], which, however, fails to remove or synthesize weather effects effectively. Although Qwen-Image-Edit [wu2025qwenimagetechnicalreport] achieves a slightly higher PickScore, it often introduces inconsistent textures and geometry. PickScore is suitable for measuring user preference, but it does not measure physical consistency and editing plausibility.
A qualitative comparison is shown in Fig. 6. In the first row, our method removes all the snowflakes and snow accumulation on the trees. Flux-Kontext fails to remove them, while Qwen-Image-Edit and WeatherWeaver mistakenly change the geometry of the scene and the car’s color. For the second row, our method removes the dense mist, recovering the color and pose of pedestrians. Flux-Kontext generates noisy textures while Qwen-Image-Edit and WeatherWeaver change the count and pose of pedestrians. In the last row, we transform the weather into a rainy day, generating natural reflection and lighting. The other methods produce unnatural lighting while struggling to preserve geometry and material. Flux-Kontext adds some rain streaks on the original image, making the sunny-day shadows on the ground look noticeably out of place. Instruct-Pix2Pix struggles to manipulate weather effects and instead performs incorrect operations resembling style transfer.
We also present the albedo and normal map obtained from our inverse renderer. Our re-rendered images align well with these maps. Weather editing in intrinsic space allows our model to completely remove weather-related artifacts, including both airborne particles and surface accumulations, while preserving geometric and material consistency. Furthermore, the disentangled material and geometry obtained from inverse rendering facilitate realistic illumination and shadow generation during re-rendering. More comparisons are shown in the supplementary.
We further compare our method with WeatherWeaver on fine-grained weather control in Fig. 7. Our editing results show natural transitions: under light to heavy rain, the road surface gradually becomes wetter; under different levels of snowfall, snow first appears along the roadside and on branches, and eventually accumulates to cover the entire road. In contrast, WeatherWeaver lacks this sense of realism and instead looks more like applying filters of increasing intensity to the original image.
User study.
We conducted a user study to evaluate the consistency and realism of different editing methods. A total of 61 participants were asked to vote on 8 cases covering weather removal and weather synthesis. The average preference for our results is 81.67%, showing that our method is preferred by users. The detailed setup and results are provided in the supplementary material.
| Method | PickScore [kirstain2023pickapicopendatasetuser] | CLIP-S | DINO-S | |||
| Sunny | Snowy | Foggy | Rainy | |||
| Rendering-based methods | ||||||
| Ours | 20.59 | 22.32 | 21.34 | 20.76 | 27.66 | 73.63 |
| RGBX | 20.40 | 19.92 | 19.71 | 20.29 | 19.00 | 55.87 |
| DiffusionRenderer | 20.24 | – | – | – | – | 43.12 |
| Pixel-space editing methods | ||||||
| Flux-Kontext | 20.72 | 22.25 | 20.99 | 19.46 | 24.46 | 85.50 |
| Qwen-Image-Edit | 20.77 | 22.43 | 21.82 | 21.56 | 27.14 | 53.70 |
| WeatherWeaver | 20.13 | 21.41 | 20.93 | 20.25 | 26.78 | 67.01 |
| Instruct-Pix2Pix | 20.27 | 21.39 | 20.98 | 19.83 | 23.79 | 22.41 |
4.2 Comparison with weather restoration methods
We present qualitative comparisons with AWRaCLe [rajagopalan2025awracle] and Histoformer [sun2024restoringimagesadverseweather] in Fig. 8. The advantage of our method is its ability to effectively remove various degradations in adverse weather images, such as airborne particles (like snowflakes), snow on the ground, and overall illumination. The weather restoration methods only remove particles while failing to change surface material or lighting conditions.
To further verify the improvement brought by different models to downstream applications, we choose object detection and semantic segmentation as tasks, evaluating performance improvement before and after weather editing. Specifically, we chose DETR [carion2020detr] and Segformer [xie2021segformer] as base models. When applying object detection and semantic segmentation models to our re-rendered images, both tasks achieve more accurate and consistent results, as shown in the lower rows of Fig. 8. Following this way, we apply weather editing and evaluate on the validation set of the ACDC benchmark [sakaridis2021acdc], and the results are shown in Tab. 2.
| mIOU | ||||
| DETR | 56.56 | 13.15 | 47.00 | – |
| DETR + ours | 61.32 | 24.60 | 54.87 | – |
| Segformer | – | – | – | 24.13 |
| Segformer + ours | – | – | – | 30.05 |
| Absolute Gain | +4.76 | +11.45 | +7.87 | +5.92 |
| Method | Albedo | Normal | Roughness | Metallicity | Irradiance | |||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | MAE | PSNR | LPIPS | PSNR | LPIPS | PSNR | LPIPS | |
| IID | 7.80 | 0.26 | 0.63 | – | – | – | 10.30 | 0.55 | 12.37 | 0.64 | – | – |
| IID (w/ finetune) | 11.55 | 0.53 | 0.40 | – | – | – | 12.34 | 0.43 | 12.22 | 0.55 | – | – |
| RGBX | 9.66 | 0.44 | 0.47 | 11.90 | 0.41 | 15.51 | 13.62 | 0.55 | – | – | 16.24 | 0.58 |
| RGBX (w/ finetune) | 11.35 | 0.59 | 0.37 | 16.14 | 0.49 | 7.05 | 13.65 | 0.57 | 11.96 | 0.66 | 16.38 | 0.69 |
| GeoWizard | – | – | – | 16.24 | 0.54 | 12.47 | – | – | – | – | – | – |
| IDArb | 6.40 | 0.48 | 0.65 | 10.77 | 0.43 | 22.42 | 10.70 | 0.62 | 14.66 | 0.62 | – | – |
| DiffusionRenderer | 11.91 | 0.64 | 0.34 | 16.43 | 0.70 | 28.68 | 11.31 | 0.42 | 10.05 | 0.43 | – | – |
| Ours | 27.99 | 0.86 | 0.35 | 25.06 | 0.84 | 4.24 | 25.81 | 0.23 | 29.29 | 0.04 | 29.66 | 0.22 |
| Ours (w/o IMAA) | 26.78 | 0.84 | 0.43 | 23.63 | 0.79 | 6.33 | 24.60 | 0.25 | 28.16 | 0.05 | 26.99 | 0.32 |
4.3 Evaluation for components
4.3.1 Inverse rendering
In this part, we first conduct quantitative evaluations on WeatherSynthetic, then we evaluate on real-world TransWeather [valanarasu2021transweather] datasets. We show the comparison between our method and existing methods on the test set of our WeatherSynthetic in Tab. 3. Our method outperforms existing approaches across all evaluation metrics. We fine-tune IID [kocsis2024intrinsic] and RGBX [zeng2024rgb] with the same training steps on our WeatherSynthetic. Their performance improves, but they fail to provide high-quality estimation. An overall qualitative comparison is shown in the supplementary.
We show a comparison of real images of heavy rain in Fig. 9. All other methods fail to give faithful estimations, while our WeatherDiffusion provides reasonable estimations. We further validate the consistency of our inverse renderer across weather conditions in the supplementary materials. For each scene, we run the inverse renderer on images captured under different weather types and compute the PSNR between the recovered intrinsic maps and those obtained under sunny conditions.
4.3.2 Forward rendering
A comparison of forward rendering between WeatherDiffusion, RGBX, and DiffusionRenderer [liang2025diffusionrenderer] is shown in Fig. 10. We use the prompt “A sunny day in the city.” for ours and RGBX, and provide an environment lighting for DiffusionRenderer. Our WeatherDiffusion produces images that better align with the text description than RGBX, and avoid abnormal textures and illuminations. DiffusionRenderer fails to recover all details, such as road signs and distant buildings. As shown in Tab. 1, our method achieves the highest metrics in rendering-based methods on all weather conditions. More results are shown in the supplementary.
4.4 Ablation study
Effect of IMAA.
We train an inverse renderer without IMAA for the same steps. As shown in Tab. 3, the model without IMAA behaves poorly than our full model. We show a qualitative result in part (a) of Fig. 11. With the attention guidance related to the map provided by IMAA, the model produced more refined geometry and material predictions and successfully identified the metallic handrail, assigning it a reasonable level of metallicity.
Effect of WeatherReal.
We explore the effect of WeatherReal. We train the forward renderer with only the synthetic dataset and with the WeatherReal dataset. The qualitative results are shown in part (b) of Fig. 11. We use the same intrinsic maps as input and evaluate each model. After training solely on the synthetic dataset, the forward renderer fails to reach high realism, resulting in unrealistic lighting and objects. After introducing WeatherReal, the model learn the distribution of the real-world data and then generates high-quality rendered images. We explore more ablation of datasets in the supplementary material.
4.5 Limitations
Our framework is designed for single-image weather editing in the intrinsic space. Temporal modeling introduces additional factors, such as object motion and occlusion changes, that are orthogonal to our core contribution and require a video prior. Therefore, the current framework does not guarantee temporal consistency in video sequences, as shown in Fig. 12. Extending WeatherDiffusion to videos would likely be possible by building on a video diffusion model, similar to DiffusionRenderer [liang2025diffusionrenderer]. However, these models typically demand substantially more training data, computational resources, and operate at lower resolutions, making them difficult to apply to driving scenes directly. Building a temporally consistent framework is an important direction for future work, but it remains beyond the scope of this paper.
5 Conclusion
We propose WeatherDiffusion, a novel framework for controllable weather editing in intrinsic space. Our approach achieves robust intrinsic decomposition across diverse weather and illumination conditions while enabling controlled weather editing based on maps and prompts. For the inverse renderer, we propose IMAA to provide attention guidance to help the model focus on semantically important regions. For the forward renderer, we leverage CLIP interpolation and diffusion priors to achieve fine-grained weather control. Last, we construct two datasets, WeatherSynthetic and WeatherReal, containing intrinsic maps to address the lack of large-scale autonomous driving rendering datasets under varied weather conditions. Our WeatherDiffusion demonstrates performing weather editing in intrinsic space can achieve controllable and plausible editing while preserving geometry and material. Future work includes extending the framework to video-based weather editing and collecting more diverse and realistic data.